21. Based on the financial statements provided below please fill the missing numbers and calculate the requested financial raNeed help on all of them please thank you!!
Answer #1

a.

Total current assets = cash and equivalents + Accounts receivable + Inventories

62,500 = Cash + 16,000+42,500

Cash = $4,000

Cash and equivalents of 2012 = $4,000

b.

Accumulated depreciation of 2011 = $28,000

Depreciation expense of 2012 = $10,000

Accumulated depreciation of 2012 = Accumulated depreciation of 2011 + Depreciation expense of 2012

= 28,000+10,000

= $38,000

c.

Total assets =Total liabilities and Owners equity

150,500 = 69,298+22,950+31,500+ Retained earnings

Retained earnings = $26,752

d.

Quick ratio = (Cash + Accounts receivable)/ Current liabilities

= (4,000+42,500)/69,298

= 0.67

Kindly comment if you need further assistance. Thanks‼!